# Fast and reliable confidence intervals for a variance component or proportion

### v1

@@ 4 Computational considerations
In this section, we consider the computational aspects of our test-statistics. Like popular competing methods, we require a one-time preprocessing step where the eigendecomposition of $K=O\Lambda O^{\mathrm{\scriptscriptstyle T}}$ is used to construct $Y_{O}=O^{\mathrm{\scriptscriptstyle T}}Y$ and $X_{O}=O^{\mathrm{\scriptscriptstyle T}}X$. In practice it is common to use the same $K$ for many different choices of response vector $Y$, with separate inference on $h^{2}$ for each response vector. Then, the pre-processing step is only performed once.
@@
Recall $V_{2}=h^{2}\Lambda+(1-h^{2})I_{n}$ and define $\tilde{P}=V_{2}^{-1/2}X_{O}(X_{O}^{\mathrm{\scriptscriptstyle T}}V_{2}^{-1}X_{O})^{-1}X_{O}^{\mathrm{\scriptscriptstyle T}}V_{2}^{-1/2}$, $\tilde{Q}=I_{n}-\tilde{P}$, $\tilde{\beta}=(X_{O}^{\top}V_{2}^{-1}X)^{-1}X_{O}^{\top}V_{2}^{-1}Y_{O}$, and $\tilde{R}=Y_{O}-X_{O}\tilde{\beta}$. Let $\tilde{R}_{i}$ denote the $i$th element of $\tilde{R}$ and $Q_{ii}$ the $i$th diagonal element of $Q$. With these definitions the score function is
@@
Both terms in (13) involve the summation of $n$ terms. Construction of $\tilde{R}$ requires $O(np)$ operations; and computation of all $\tilde{Q}_{ii}$ requires only $O(np^{2})$ operations. We do not compute or store the entire matrix $\tilde{Q}$.
To compute the information efficiently, note $Q=O\tilde{Q}O^{\top}$, so by Proposition 1,
@@
We have created an R package lmmvar for computing our test-statistics and confidence intervals, available for download at github.com/yqzhang5972/lmmvar.
Our confidence interval for $h^{2}$ can be obtained by evaluating $T_{n}^{R}(h^{2})$ from at equally spaced points in $[0,1)$ and comparing the values to quantiles from the chi-square distribution with one degree of freedom. In practice, however, it is typically possible, and faster, to do a search for the points where $T_{n}^{R}(h^{2})$ crosses the desired quantile. More specifically, when $T_{n}^{R}(h^{2})$ is strictly quasiconvex in $h^{2}$, then $\mathcal{R}_{n}(\alpha)$ is an interval, so we need only search for points $a$ such that $T_{n}^{R}(a)=q_{1,1-\alpha}$. While we have not been able to prove $T_{n}^{R}(h^{2})$ is quasiconvex in general due to technical difficulties, we have not encountered an example where it is not. Typically, $T_{n}^{R}(h^{2})$ looks roughly like one of the three cases in Fig. 2. Thus, in our software we first find a point $a^{\prime}$ such that $T_{n}^{R}(a^{\prime})<q_{1,1-\alpha}$. More specifically, we start a ternary search for the minimum of $T_{n}^{R}$ and terminate the first time a negative value is encountered. Then, under quasiconvexity, the intervals $[0,T_{n}^{R}(a^{\prime})]$ and $[T_{n}^{R}(a^{\prime}),1)$ contain lower and upper bounds of the interval $\mathcal{R}_{n}(\alpha)$, respectively. We then use a bisection search to find $b\in[0,T_{n}^{R}(a^{\prime})]$ such that $T_{n}^{R}(b)=q_{1,1-\alpha}$, and similarly for $[T_{n}^{R}(a^{\prime}),1)$. This procedure often requires less computing time than an exhaustive grid search.
@@
Figure 2: The value $T_{n}^{R}(h^{2})$ with $n=1000$, data generated under the same model as in Fig. 1, and the true $h^{2}$, here denoted $h_{*}^{2}$, taking values $\{0.0001,0.5,0.999\}.$ Horizontal red lines denote the $95$th percentile of the $\chi^{2}_{1}$ distribution.
@@ 5.1 Finite-sample interval width and coverage probabilities
In each setting considered, we generate $X\in\mathbb{R}^{n\times 5}$ with independent and identically distributed standard normal entries, and then generate $Y\sim\mathcal{N}_{n}(X\beta,h^{2}K+(1-h^{2})I_{n})$ (i.e., $\sigma^{2}=1$) with $\beta=0$. For every combination of the triplet $(n,h^{2},\rho)$, we generate $10^{4}$ independent realizations of $Y$ and compute the corresponding 95% confidence intervals for $h^{2}$.
@@
The top panel of Fig. 3 displays interval widths for different values of $\rho$ and $h^{2}$. When $\rho=0.1$, $K\approx I_{n}$, which means $h^{2}$ is approximately unidentifable. Consequently, interval widths are too large to be practically useful, despite having the nominal coverage level. With $\rho=0.5$ or $0.95$, the widths are reasonably small and decrease as $n$ increases. When $\rho=0.999$, the eigenvalues of $K$ are approximately $(n,0,\dots,0)$, which appear to lead to longer confidence intervals. Nevertheless, as $n$ increases, interval widths decrease as expected. As $n$ approaches 2000, the widths are small enough to be useful in practice.
@@ 5.2 Comparison to simulation-based test-statistics
Existing methods for constructing confidence intervals for $h^{2}$ often rely on simulation. For example, Schweiger et al., (2016) uses the parametric bootstrap to approximate the distribution of the REML estimator of $h^{2}.$ In comparison to a naive implementation of the parametric bootstrap, their implementation is very efficient. Specifically, to estimate the distribution of the REML estimator when the true heritability is $h^{2}$, their method requires (i) drawing $N$ independent samples from (1) with $\sigma^{2}=1$ and null hypothesis value $h^{2}$ given; and (ii) for each sample, finding an interval containing a local maximizer of the restricted log-likelihood. Since each sample is needed for step (ii), Schweiger et al., (2016) show that it is sufficient to draw samples from $\mathcal{N}_{n}(0,I_{n})$ in step (i) and modify step (ii) accordingly. The modified step (ii) is performed by evaluating the derivative of the restricted log-likelihood at fixed grid points. However, for a given $h_{2}$ the derivative must, in general, be evaluated at many grid points; and this procedure must be repeated for all $N$ samples. Constructing confidence intervals thus requires repeating this procedure many times, which can be computationally burdensome.
@@
In follow-up work, Schweiger et al., (2018) proposed a stochastic approximation to the confidence interval from Schweiger et al., (2016). This approximation estimates the upper and lower bounds of the confidence interval using the Robbins–Monro algorithm. Like our test-statistic, after the same preprocessing step, the operations required for constructing the intervals from Schweiger et al., (2016) and Schweiger et al., (2018) scale linearly with $n$.
@@
In Fig. 4, we compare the time needed to compute our interval to that of Schweiger et al., (ALBI, 2016) and Schweiger et al., (FIESTA, 2018). The figure indicates our confidence intervals can be computed most efficiently of the three methods. For example, then $n=100$ and $h^{2}=0.01$ our method took on average $3.7\times 10^{-4}$ seconds compared to 3.11 seconds and 45.38 seconds for FIESTA and ALBI, respectively. When $n=3000$ and $h^{2}=0.01$, our method took 0.07 seconds on average, whereas FIESTA and ALBI took 511.18 seconds and 1908.68 seconds, respectively. ALBI and FIESTA are written in Python, while our software is written in a combination of R and C++. All timings were obtained simultaneously on the University of Florida’s high-performance computing cluster, HiPerGator.
We also compared the width of our confidence intervals to those of Schweiger et al., (2018). We found that under a broad range of settings, the distribution of our interval’s width was similar to that of Schweiger et al., (2018). In our simulations with $n=1000$, we recorded the standardized difference in widths (defined as Schweiger et al., (2018)’s width minus our interval’s width, divided by our interval’s width). On average the standardized differences were $0.24,-0.03$, and $0.005$ when $h^{2}$ was $0.01,0.1$, and $0.5$, respectively. Median standardized differences were $-0.07,-0.03$, and $0.005$.
@@
Figure 4: The average computing times for constructing one confidence interval for $h^{2}$ using our test-statistic $T_{n}^{R}$ (”Proposed”), the simulation-based interval construction method from Schweiger et al., (2016) (”ABLI”), and the stochastic approximation of the interval from Schweiger et al., (2016) proposed by Schweiger et al., (2018) (”FIESTA”). The shaded regions are 95% confidence bands. All times exclude the required eigendecomposition of $K$. Data were generated under the same model as in Fig. 1.
@@
Figure 5: (Left panels) Two-sided $95\%$ confidence intervals for $h^{2}$ in anterior (top row) and posterior (bottom row) regions of a mouse’s brain tissue, sorted according to their lower bound. (Middle panels) Histograms of the lower bounds of $95\%$ one-sided confidence intervals for $h^{2}$ in both anterior and poster regions. (Right panel) Expression counts of SNAP25, one of the most spatially variable genes, across the posterior region. We display counts, but our analysis is performed on log-normalized expression.
@@ 5.3 Gene ranking in spatially-resolved transcriptomics
In Section 1, we described one application of our method in the analysis of spatially-resolved transcriptomic data. The rightmost panels of Fig. 5 display data generated in a spatially-resolved RNA sequencing experiment performed on a mouse brain tissue sample slice. In such an experiment, one measures the expression of thousands of genes (the number of RNA reads aligning with each gene) from cells in 50 micron diameter spatial locations (known as spots) across the tissue sample slice. These locations are represented by the colored circles in the rightmost panels of Fig. 5.
In this context, for one gene, $h^{2}$ is the fraction of variability in expression attributed to the spatial orientation of the tissue sample. In a typical analysis, one tests $h^{2}$ for more than 10,000 genes. Specially, after normalization, it is common to assume that for the $\ell$th gene,
$Y_{\ell}\sim\mathcal{N}_{n}(X\beta,\sigma^{2}\{h_{\ell}^{2}K+(1-h_{\ell}^{2})I_{n}\}),\quad\ell\in\{1,\dots,d\},$ (14)
@@
To demonstrate the usefulness of our method in this application, we analyzed two slices of a mouse’s brain tissue (anterior and posterior regions), which are available as the stxbrain data in the SeuratData R package. These data were generated on the 10X Visium platform. Assuming (14) with $a=0.02$ (chosen based on results from Weber et al., (2022)), after quality control of filtering out low-expressed genes and boundary spots in the tissues, $n=2380$ for $d=15117$ genes in the anterior slice and $n=3011$ for $d=14738$ genes in the posterior slice.

### v2

@@ 4 Computational considerations
In this section, we consider the computational aspects of our test-statistics. Like popular competing methods, we require a one-time preprocessing step where the eigendecomposition of $K=O\Lambda O^{\mathrm{\scriptscriptstyle T}}$ is used to construct $Y_{O}=O^{\mathrm{\scriptscriptstyle T}}Y$ and $X_{O}=O^{\mathrm{\scriptscriptstyle T}}X$. In some applications, the same $K$ is used for many different choices of response vector $Y$, with separate inference on $h^{2}$ for each response vector. Then, the pre-processing step is only performed once.
@@
Recall $V_{2}=h^{2}\Lambda+(1-h^{2})I_{n}$ and define $\tilde{P}=V_{2}^{-1/2}X_{O}(X_{O}^{\mathrm{\scriptscriptstyle T}}V_{2}^{-1}X_{O})^{-1}X_{O}^{\mathrm{\scriptscriptstyle T}}V_{2}^{-1/2}$, $\tilde{Q}=I_{n}-\tilde{P}$, $\tilde{\beta}=(X_{O}^{\mathrm{\scriptscriptstyle T}}V_{2}^{-1}X_{O})^{-1}X_{O}^{\mathrm{\scriptscriptstyle T}}V_{2}^{-1}Y_{O}$, and $\tilde{R}=Y_{O}-X_{O}\tilde{\beta}$. Let $\tilde{R}_{i}$ denote the $i$th element of $\tilde{R}$ and $Q_{ii}$ the $i$th diagonal element of $Q$. The score function is then
@@
Both terms in (13) involve the summation of $n$ terms. Construction of $\tilde{R}$ requires $O(np)$ operations, and computation of all $\tilde{Q}_{ii}$ requires $O(np^{2})$ operations. We do not compute or store the entire matrix $\tilde{Q}$.
To compute the information efficiently, note $Q=O\tilde{Q}O^{\mathrm{\scriptscriptstyle T}}$, so by Proposition 1,
@@
We have created an R package lmmvar for computing our test-statistics and confidence intervals, available for download at github.com/yqzhang5972/lmmvar. The package also provides an implementation of the method by Crainiceanu and Ruppert, (2004).
A confidence interval for $h^{2}$ can be obtained by evaluating $T_{n}^{R}(h^{2})$ from at equally spaced points in $[0,1)$ and comparing the values to quantiles from the chi-square distribution with one degree of freedom. In practice, however, it is typically possible, and faster, to do a search for the points where $T_{n}^{R}(h^{2})$ crosses the desired quantile. More specifically, when $T_{n}^{R}(h^{2})$ is strictly quasiconvex in $h^{2}$, then $\mathcal{R}_{n}(\alpha)$ is an interval, so we need only search for points $a$ such that $T_{n}^{R}(a)=q_{1,1-\alpha}$. While we have not been able to prove $T_{n}^{R}(h^{2})$ is quasiconvex in general due to technical difficulties, we have not encountered an example where it is not. Typically, $T_{n}^{R}(h^{2})$ looks roughly like one of the three cases in Fig. 2. Thus, in our software we first find a point $a^{\prime}$ such that $T_{n}^{R}(a^{\prime})<q_{1,1-\alpha}$. More specifically, we start a ternary search for the minimum of $T_{n}^{R}$ and terminate the first time a value less than $q_{1,1-\alpha}$ is encountered. Then, under quasiconvexity and assuming $a^{\prime}$ is not on the boundary, the intervals $[0,T_{n}^{R}(a^{\prime})]$ and $[T_{n}^{R}(a^{\prime}),1)$ contain lower and upper bounds of the interval $\mathcal{R}_{n}(\alpha)$, respectively. We then use a bisection search to find $b\in[0,T_{n}^{R}(a^{\prime})]$ such that $T_{n}^{R}(b)=q_{1,1-\alpha}$, and similarly for $[T_{n}^{R}(a^{\prime}),1)$. If the point $a^{\prime}$ is on the boundary, as in the first and third plot of Fig. 2, only one bisection search is needed. In either case, this procedure often requires less computing time than a grid search.
@@
Figure 2: The value $T_{n}^{R}(h^{2})$ with $n=1000$, data generated under the same model as in Fig. 1, and the true $h^{2}$, here denoted $h_{*}^{2}$, taking values $0.0001,0.5$, and $0.999$ Horizontal red lines indicate the $95$th percentile of the $\chi^{2}_{1}$ distribution.
@@ 5.1 Finite-sample interval width and coverage probabilities
In each setting considered, we generate $X\in\mathbb{R}^{n\times 5}$ with independent and identically distributed standard normal entries, and then generate $Y\sim\mathcal{N}_{n}(X\beta,h^{2}K+(1-h^{2})I_{n})$ (i.e., $\sigma^{2}=1$) with $\beta=0$. For every combination of the triplet $(n,h^{2},\rho)$, we generate $10^{4}$ independent realizations of $Y$ and compute the corresponding 95% confidence intervals for $h^{2}$ for each realization. Thus, coverage probability and average width intervals in Figure. 3 are estimated based on $10^{4}$ independent replicates.
@@
The top panel of Fig. 3 displays interval widths for different values of $\rho$ and $h^{2}$. When $\rho=0.1$, $K\approx I_{n}$, which means $h^{2}$ is approximately unidentifiable. Consequently, interval widths are too large to be practically useful, despite having the nominal coverage level. With $\rho=0.5$ or $0.95$, the widths are reasonably small and decrease as $n$ increases. When $\rho=0.999$, the eigenvalues of $K$ are approximately $(n,0,\dots,0)$, which appear to lead to longer confidence intervals. Nevertheless, as $n$ increases, interval widths decrease as expected. As $n$ approaches 2000, the widths are small enough to be useful in practice.
@@ 5.2 Comparison to simulation-based test-statistics
Existing methods for constructing confidence intervals for $h^{2}$ often rely on simulation. For example, Schweiger et al., (2016) uses the parametric bootstrap to approximate the distribution of the restricted maximum likelihood estimator of $h^{2}.$ In comparison to a naive implementation of the parametric bootstrap, their implementation is very efficient. Specifically, to estimate the distribution of the estimator when the true heritability is $h^{2}$, their method requires (i) drawing $N$ independent samples from (2) with $\sigma^{2}=1$ and null hypothesis value $h^{2}$ given; and (ii) for each sample, finding an interval containing a local maximizer of the restricted log-likelihood. Since each sample is needed for step (ii), Schweiger et al., (2016) show that it is sufficient to draw samples from $\mathcal{N}_{n}(0,I_{n})$ in step (i) and modify step (ii) accordingly. The modified step (ii) is performed by evaluating the derivative of the restricted log-likelihood at fixed grid points. However, for a given $h^{2}$ the derivative must, in general, be evaluated at many grid points; and this procedure must be repeated for all $N$ samples. Constructing confidence intervals thus requires repeating this procedure many times, which can be computationally burdensome. Nevertheless, the procedure is faster than previous state-of-the-art simulation-based method due to Crainiceanu and Ruppert, (2004). The latter method provides an interval for $\tau=\sigma^{2}_{g}/\sigma^{2}_{e}$, which is straightforward to turn into an interval for $h^{2}=\tau/(1+\tau)$.
@@
In follow-up work, Schweiger et al., (2018) proposed a stochastic approximation to the confidence intervals of Schweiger et al., (2016). The approximation estimates the upper and lower bounds of the confidence interval using the Robbins–Monro algorithm. Like our test-statistic, after the same preprocessing step, the number of operations required for constructing their intervals scales linearly with $n$.
@@
In Fig. 4, we compare the time needed to compute our interval to those of Schweiger et al., (ALBI, 2016), Schweiger et al., (FIESTA, 2018) and Crainiceanu and Ruppert, (RLRT, 2004). The figure indicates our confidence intervals can be computed most efficiently. For example, when $n=200$ and $h^{2}=0.01$ our method took on average $7.3\times 10^{-4}$ seconds compared to 2.89 seconds, 69.98 seconds and 327.10 for FIESTA, ALBI and RLRT, respectively. When $n=2000$ and $h^{2}=0.01$, our method took 0.02 seconds on average, whereas FIESTA, ALBI and RLRT took 119.38 seconds, 736.51 seconds and 3983.35 seconds respectively. ALBI and FIESTA are written in Python, while our software and RLRT are written in R and C++. All timings were obtained simultaneously on the University of Florida’s high-performance computing cluster, HiPerGator.
We also compared the widths of our confidence intervals to those of Schweiger et al., (2016). We found that under a broad range of settings, the distribution of the widths were similar. In the Supplementary Material, we provide plots comparing the two intervals’ widths for $n\in\{200,500,1000,2000\}$ and $h^{2}$ varying. Due to the long computing times for the interval of Crainiceanu and Ruppert, (2004), we excluded it from these comparisons.
@@
Figure 4: Average computing times for one confidence interval for $h^{2}$ using the methods proposed here (“Proposed”), by Schweiger et al., (2016) (“ALBI”), by Schweiger et al., (2018) (“FIESTA”), and by Crainiceanu and Ruppert, (2004) (“RLRT”). There are shaded regions providing 95% confidence bands. All times exclude the required eigendecomposition of $K$. Data were generated under the same model as in Fig. 1.
@@
Figure 5: (Left panels) Two-sided $95\%$ confidence intervals for $h^{2}$ in anterior (top row) and posterior (bottom row) regions of a mouse’s brain tissue, sorted according to their lower bound. (Middle panels) Histograms of the lower bounds of $95\%$ one-sided confidence intervals for $h^{2}$ in both anterior and posterior regions. (Right panel) Expression counts of SNAP25, one of the most spatially variable genes, across the posterior region. We display counts, but our analysis is performed on log-normalized expression.
@@ 5.3 Gene ranking in spatially-resolved transcriptomics
In Section 1, we described one application of our method in the analysis of spatially resolved transcriptomic data. The rightmost panels of Fig. 5 display data generated in a spatially resolved RNA sequencing experiment performed on a mouse brain tissue sample. In such an experiment, one measures the expression of thousands of genes (the number of RNA reads aligning with each gene) from cells in 50 micron diameter spatial locations (known as spots) across the tissue sample slice. These spots are represented by the colored circles in the rightmost panels of Fig. 5; expression is measured separately in each spot.
In this context, for one gene, $h^{2}$ is the fraction of variability in expression attributed to the spatial orientation of the tissue sample. In a typical analysis, one tests $h^{2}$ for more than 10,000 genes. Specifically, it is common to assume that for the $\ell$th gene, the vector of normalized random expression in all spots $Y_{\ell}$, satisfies
$Y_{\ell}\sim\mathcal{N}_{n}(X\beta_{\ell},\sigma_{\ell}^{2}\{h_{\ell}^{2}K+(1-h_{\ell}^{2})I_{n}\}),\quad\ell\in\{1,\dots,d\},$ (14)
@@
To demonstrate the usefulness of our method in this application, we analyzed two separate slices of a mouse’s brain tissue (anterior and posterior regions), which are available as the stxbrain data in the SeuratData R package. These data were generated on the 10X Visium platform. Assuming (14) with $a=0.02$ (chosen based on results from Weber et al., (2022)), after quality control of filtering out low-expressed genes and boundary spots in the tissues, $n=2380$ for $d=15117$ genes in the anterior slice and $n=3011$ for $d=14738$ genes in the posterior slice.